THE INTERSLELLAR OH AND H_{2}O MASERS

\underline{HANPING LIU} and JIN SUN, Department of Physics, Beijing Normal University, Beijing 100875.

The masers in the regions of star formation are mainly both H_{2}O and OH masers. They are related to the early evolution of stars. That is why an understanding of the excitated mechanism is of central importance for the discovery of physical characteristic of star formation regions. Unfortunately, the excitation mechanism is not understood well yet. So far there are two kinds of known mechanisms: Collision and Radiation mechanism. But there are still serious puzzles with these mechanisms. The criticisms of radiation excitation mechanism were its short of pumping photons and its dissociation of H_{2}O and OH molecules by the radiation itself. We notice that the remarkable features of star-formation regions are its ultra-cold environment, its strong VUV radiation and its coicident appearence of the H_{2}O and OH masers. Therefore, we argue a new excitation mechanism. It can unified interpret both H_{2}O and OH interstellar masers. This mechanism uses a wider frequency range of photons and includes a scheme for the regeneration of interstellar H_{2}O and OH molecules. Those are different from former radiation excitation mechanism.